### v1

@@ Acknowledgments
Grade 1 371 (45.4) 217 (50.9) 103 (43.6) 149 (52.3) 94 (48.0) 63 (48.5) 91 (50.8) 78 (54.5)
Grade 2 284 (34.8) 122 (28.6) 89 (37.7) 97 (34.0) 71 (36.2) 39 (30.0) 57 (31.8) 44 (30.8)
Grade 3 162 (19.8) 87 (20.4) 44 (18.6) 39 (13.7) 31 (15.8) 28 (21.5) 31 (17.3) 21 (14.7)
@@
Academic Medical Centers 318 (38.9) 168 (39.4) 113 (47.9) 104 (36.5) 53 (27.0) 71 (54.6) 68 (38.0) 48 (33.6)
Community 336 (41.1) 124 (29.1) 68 (28.8) 106 (37.2) 69 (35.2) 27 (20.8) 44 (24.6) 48 (33.6)
Specialty 16 ( 2.0) 5 ( 1.2) 1 ( 0.4) 2 ( 0.7) 2 ( 1.0) 3 ( 2.3) 1 ( 0.6) 1 ( 0.7)
Teaching 132 (16.2) 110 (25.8) 54 (22.9) 61 (21.4) 66 (33.7) 29 (22.3) 60 (33.5) 44 (30.8)
@@
Small (<100) 21 ( 2.6) 7 ( 1.6) 3 ( 1.3) 11 ( 3.9) 2 ( 1.0) 1 ( 0.8) 13 ( 7.3) 3 ( 2.1)
Medium (100-299) 213 (26.1) 71 (16.7) 47 (19.9) 72 (25.3) 32 (16.3) 22 (16.9) 36 (20.1) 33 (23.1)
Large (300+) 583 (71.4) 348 (81.7) 186 (78.8) 202 (70.9) 162 (82.7) 107 (82.3) 130 (72.6) 107 (74.8)
@@
Family/Internal medicine 427 (52.3) 192 (45.1) 138 (58.5) 105 (36.8) 81 (41.3) 80 (61.5) 43 (24.0) 52 (36.4)
Hematology 13 ( 1.6) 3 ( 0.7) 3 ( 1.3) 3 ( 1.1) 0 ( 0.0) 0 ( 0.0) 1 ( 0.6) 1 ( 0.7)
Gynecology and obstetrics 155 (19.0) 83 (19.5) 39 (16.5) 71 (24.9) 48 (24.5) 13 (10.0) 46 (25.7) 27 (18.9)
Oncology 105 (12.9) 63 (14.8) 28 (11.9) 39 (13.7) 31 (15.8) 18 (13.8) 35 (19.6) 29 (20.3)
Radiology 25 ( 3.1) 10 ( 2.3) 4 ( 1.7) 5 ( 1.8) 5 ( 2.6) 3 ( 2.3) 7 ( 3.9) 5 ( 3.5)
Other specialty 15 ( 1.8) 7 ( 1.6) 6 ( 2.5) 8 ( 2.8) 4 ( 2.0) 4 ( 3.1) 5 ( 2.8) 3 ( 2.1)
Unknown 77 ( 9.4) 68 (16.0) 18 ( 7.6) 54 (18.9) 27 (13.8) 12 ( 9.2) 42 (23.5) 26 (18.2)
Supporting Materials
Appendix A Multivariate Bernoulli Mixture Model
Figure A.1 shows the pairwise correlations of the selected NSDoH variables. Given that we used a small geographic unit (i.e., census tracts), some variables are highly correlated. In the case of the MBMM, the variance-covariance matrix of a mixture component of independent Bernoulli distributions is not diagonal and does not assume independence between variables within clusters; therefore, this model can accommodate neighborhood-level data.
Figure A.1: Pairwise Pearson’s Correlations of NSDoH variables from ACS 2015-2019.
We assumed that the observed data comes from a mixture of $K$ independent Bernoulli distributions. Here, $K$ is the number of mixture components (i.e., clusters) [^1] In the manuscript text we refer to clusters as profiles. included in the model, while the true number of clusters in the data is $K_{0}<K$ and is determined by model estimation. That is, we assume that for a binary data matrix $\mathbf{X}=\{\mathbf{x_{1}},...\mathbf{x_{n}}\}$, each $\mathbf{x_{i}}=\{x_{i,1},...,x_{i,p}\}$ is such that $p(x_{i})=\sum_{k=1}^{K}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}$ as described in the methods section 2.2, where $\boldsymbol{\pi}=\{\pi_{1},...,\pi_{K}\}$ is the probability vector for the cluster assignments (i.e., the probability that a census tract belongs to cluster $k\in\{1,...,K\}$) and $\sum_{k=1}^{K}\pi_{k}=1$, and the probability matrix $\boldsymbol{\theta}=\{\theta_{j|k}\}^{p\times K}$ represents the probability of a high level of exposure to NSDoH variable $j$ for a census tract given assignment to cluster $k$. Therefore, the observed likelihood for the specified model is
@@ Appendix A Multivariate Bernoulli Mixture Model
We assume no prior knowledge for the initialization of the parameters. Consequently, all parameters are initialized with non-informative priors detailed below. The true number of NSDoH profiles is unknown. Using the upper bound $K_{max}$( also referred to as $K$ in our model description above), we impose an overfitted finite mixture model[44]. We fit the model with a large upper bound on the number of clusters, $K=50$, coupled with a prior on the number of clusters $K<K_{max}$, where $K$ is treated by the model as another unknown parameter, and allow a data-driven approach to estimating $K$[44, 24]. We impose the following priors on model parameters:
@@
$\boldsymbol{\pi}\mid K\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{K_{m}ax}),\quad\text{where }\gamma_{k}=1\;\forall\;k.$
@@
Additionally, we conducted sensitivity analyses using other priors allowed in the software package. Namely, we used $K\sim Uniform(1,K_{max})$, and $\boldsymbol{\pi}|K\sim\text{Dirichlet}(\frac{1}{K_{max}},...,\frac{1}{K_{max}})$, but we found no significant differences, in fact, the model behaved better with the priors specified above. The following full conditional distributions were used to update the model parameters:
@@
Given that they are smaller geographic units, some census tracts have missing information on some of the NSDoH variables (ranging from 14-24 census tracts with missing across all variables). This model can handle missing data by imputing these values using the parameter’s posterior mean estimates. The $MC^{3}$ for posterior computation was run for 15,000 iterations, with thinning every 10 iterations (i.e., retaining every 10th sample) and the first 5000 iterations removed as part of posterior samples post-processing. For the MCMC chain heating parameter denoted as $h_{m},m=2,...,M$, where $M$ is the total number of parallel chains, we used incremental heating where the heat of the $m$’th chain is $h_{m}=1/[1+\Delta T\times(m-1)]$, and we tuned the parameter $\Delta T$ such that swaps between chains were accepted 20%- 60% of time [55]. Through this process, we found that four (4) heated chains for the $MC^{3}$ algorithm produced good mixing, and for this type of data, smaller $\Delta T=0.025$ works better. Posterior mean estimates were calculated from the remaining 1000 iterations collected from the sampler’s output. The generated MCMC samples were postprocessed using the Equivalence Classes Representatives (ECR) algorithm to overcome label-switching identifiability issues inherent in Bayesian mixture models[56]. The most probable number of clusters ($K_{map}$) given the data is inferred, and accordingly, the NSDoH profile assignment probabilities for each census tract are estimated after reordering with the ECR algorithm given $K_{map}$. Figure A.2 shows the distribution of the assignment probabilities for all NSDoH profiles given $K_{map}$ from the model.
The model yielded nine (9) NSDoH profiles. However, we applied additional post-processing steps to reclassify census tracts that were part of smaller NSDoH profiles, defined as profiles containing $<$ 5% of the total census tracts. This step was taken to produce more meaningful and interpretable clusters. Reclassification was based on the second-highest posterior probability, as it was not significantly lower than the highest probability. Additionally, as shown in Figure A.2, some cluster assignment probabilities were low, which may suggest that there is unobserved heterogeneity (e.g., due to unobserved NSDoH variables), which the model did not account for. The final clusters were qualitatively described based on thematic domains to create NSDoH profiles as discussed in Section 3.1.
Figure A.2: Distribution of assignment probabilities for each NSDoH profile.
@@
Table A.1 shows the distribution of age and race/ethnicity within the census tracts used to further characterize NSDoH profiles in Section 3.2.
@@
Table A.1: Distribution of population characteristics within census tracts in a given NSDoH Profile.
@@
Profile 1 Profile 2 Profile 3 Profile 4 Profile 5 Profile 6 Profile 7 Profile 8
Mean (SD) n = 401 n = 378 n= 185 n = 142 n= 109 n = 100 n = 87 n = 76
Census tract population 5090 (1860) 4220 (1690) 4300 (1730) 5000 (1820) 4910 (1870) 4700 (1650) 4140 (1560) 4550 (1990)
Median Age 44.7 (6.24) 35.3 (5.23) 35.3 (6.76) 46.2 (4.72) 42.1 (5.07) 34.5 (5.13) 46.6 (8.33) 41.6 (7.02)
Ages 20-24 5.36 (5.26) 7.96 (4.29) 10.4 (9.39) 5.47 (1.67) 7.12 (6.28) 9.92 (8.06) 6.45 (6.35) 7.32 (8.07)
Ages 25-34 8.94 (3.57) 16.4 (4.37) 24.4 (12.3) 10.7 (2.51) 12.4 (3.57) 25.4 (7.94) 10.6 (3.97) 14.2 (4.55)
Ages 35-44 11.6 (2.89) 13.1 (3.25) 12.6 (3.75) 11.2 (2.33) 12.3 (3.19) 13.1 (3.87) 10.3 (3.03) 12.1 (3.72)
Age 45-54 15.5 (3.21) 12.6 (3.10) 10.9 (4.45) 15.1 (2.52) 14.2 (3.23) 10.7 (3.56) 13.2 (3.45) 12.9 (3.48)
Age 55-59 8.07 (2.05) 6.09 (1.84) 5.30 (2.52) 8.52 (1.85) 7.91 (1.98) 5.60 (2.39) 7.84 (2.49) 7.65 (3.65)
Ages 60-64 7.21 (2.31) 5.44 (2.06) 4.86 (2.29) 8.16 (1.79) 7.21 (2.20) 5.03 (2.12) 8.23 (2.79) 7.05 (2.60)
Ages 65-74 11.0 (4.80) 7.17 (2.61) 7.51 (3.40) 11.8 (3.34) 9.50 (2.55) 7.21 (2.97) 13.2 (5.37) 9.83 (3.66)
Ages 75-84 5.43 (2.32) 3.60 (1.93) 3.78 (2.23) 5.48 (2.61) 4.87 (2.02) 3.73 (1.83) 6.47 (3.08) 5.02 (2.64)
Age 85 or older 2.60 (1.81) 1.78 (1.31) 1.99 (1.91) 2.43 (1.81) 2.56 (1.72) 2.01 (1.75) 2.98 (1.74) 2.69 (2.44)
Non-Hispanic White 85.9 (8.81) 41.9 (24.8) 72.9 (12.0) 92.2 (4.52) 71.8 (18.8) 57.4 (12.9) 86.7 (9.43) 81.6 (15.1)
Non-Hispanic Black 2.18 (3.54) 16.1 (18.9) 4.36 (4.09) 1.29 (1.40) 9.19 (12.0) 10.1 (8.27) 3.35 (4.20) 4.73 (7.07)
Non-Hispanic Asian 6.26 (6.32) 5.79 (9.00) 12.1 (8.87) 1.41 (1.81) 3.81 (4.16) 15.0 (10.1) 2.51 (2.53) 1.65 (2.44)
Hispanic or Latino 3.40 (2.48) 31.7 (22.4) 7.17 (4.00) 3.08 (2.91) 11.6 (9.70) 13.9 (10.5) 4.71 (4.40) 8.91 (12.4)
Appendix B Regression Analysis Additional
The main outcome is optimal care, defined as adherence to National Comprehensive Cancer Network (NCCN) guidelines. These guidelines recommend a combination of therapies depending on the stage and grade of the tumor (Figure B.1). We used year-specific guidelines spanning the study period due to gradual changes in NCCN treatment guidelines. We determined optimal care by comparing the treatment received and the treatment recommended by NCCN. For example, we know that surgery is the first course of treatment, and thus if a woman received surgery alone or surgery and other additional therapies, as shown in Figure B.1, based on the stage and grade of their tumor, then she is classified as having received optimal care. We combined all the corresponding data to create a binary outcome variable, where:

### v2

@@ Acknowledgments
Grade 1 442 (46.4) 229 (52.2) 253 (51.2) 155 (45.6) 87 (46.5)
Grade 2 340 (35.7) 119 (27.1) 158 (32.0) 125 (36.8) 61 (32.6)
Grade 3 170 (17.9) 91 (20.7) 83 (16.8) 60 (17.6) 39 (20.9)
@@
Academic Medical Centers 347 (36.4) 165 (37.6) 163 (33.0) 168 (49.4) 100 (53.5)
Community 395 (41.5) 125 (28.5) 161 (32.6) 96 (28.2) 45 (24.1)
Specialty 16 ( 1.7) 6 ( 1.4) 3 ( 0.6) 3 ( 0.9) 3 ( 1.6)
Teaching 168 (17.6) 126 (28.7) 151 (30.6) 72 (21.2) 39 (20.9)
@@
Small (<100) 35 ( 3.7) 7 ( 1.6) 15 ( 3.0) 2 ( 0.6) 2 ( 1.1)
Medium (100-299) 257 (27.0) 67 (15.3) 97 (19.6) 73 (21.5) 32 (17.1)
Large (300+) 660 (69.3) 365 (83.1) 382 (77.3) 265 (77.9) 153 (81.8)
@@
Family/Internal medicine 451 (47.4) 198 (45.1) 170 (34.4) 193 (56.8) 106 (56.7)
Hematology 13 ( 1.4) 2 ( 0.5) 2 ( 0.4) 5 ( 1.5) 2 ( 1.1)
Gynecology and obstetrics 201 (21.1) 88 (20.0) 115 (23.3) 62 (18.2) 16 ( 8.6)
Oncology 136 (14.3) 69 (15.7) 84 (17.0) 35 (10.3) 24 (12.8)
Radiology 28 ( 2.9) 9 ( 2.1) 16 ( 3.2) 7 ( 2.1) 4 ( 2.1)
Other specialty 18 ( 1.9) 5 ( 1.1) 11 ( 2.2) 10 ( 2.9) 8 ( 4.3)
Unknown 105 (11.0) 68 (15.5) 96 (19.4) 28 ( 8.2) 27 (14.4)
Web-based supplemental materials for “A Bayesian Mixture Model Approach to Examining Neighborhood Social Determinants of Health Disparities in Endometrial Cancer Care in Massachusetts”
Carmen B Rodríguez, Stephanie M Wu, Stephanie Alimena, Alecia J McGregor and Briana JK Stephenson
1 Multivariate Bernoulli Mixture Model
Figure 1.1 shows the distribution of all the NSDoH variables. The distribution of these variables is skewed; therefore, as indicated by the vertical red line, we chose to dichotomize all variables for the MBMM model based on the median. We use the median because it is a robust non-parametric measure of central tendency unaffected by outliers. It ensures an equal data split, creating two comparison groups balanced in sample size.
Figure 1.1: Distributions of NSDoH variables from ACS 2015-2019 5-year estimates for Massachusetts.
Figure 1.2 shows the pairwise correlations of the selected NSDoH variables. Given that we used a small geographic unit (i.e., census tracts), some variables are highly correlated. In the case of the MBMM, the variance-covariance matrix of a mixture component of independent Bernoulli distributions is not diagonal and does not assume independence between variables within clusters; therefore, this model can accommodate neighborhood-level data.
Figure 1.2: Pairwise Pearson’s Correlations of NSDoH variables from ACS 2015-2019.
We assumed the observed data comes from a mixture of $K$ independent Bernoulli distributions. Here, $K$ is the number of mixture components (i.e., clusters) [^1] In the manuscript text, we refer to clusters as profiles. included in the model, while the true number of clusters in the data is $K_{0}<K$ and is determined by model estimation. That is, we assume that for a binary data matrix $\mathbf{X}=\{\mathbf{x_{1}},...\mathbf{x_{n}}\}$, each $\mathbf{x_{i}}=\{x_{i,1},...,x_{i,p}\}$ is such that $p(x_{i})=\sum_{k=1}^{K}\pi_{k}\prod_{j=1}^{p}\theta_{j|k}^{x_{i,j}}(1-\theta_{j|k})^{1-x_{i,j}}$ as described in the methods Section 2.2, where $\boldsymbol{\pi}=\{\pi_{1},...,\pi_{K}\}$ is the probability vector for the cluster assignments (i.e., the probability that a census tract belongs to cluster $k\in\{1,...,K\}$) and $\sum_{k=1}^{K}\pi_{k}=1$, and the probability matrix $\boldsymbol{\theta}=\{\theta_{j|k}\}^{p\times K}$ represents the probability of a high level of exposure to NSDoH variable $j$ for a census tract given assignment to cluster $k$. Therefore, the observed likelihood for the specified model is
@@ 1 Multivariate Bernoulli Mixture Model
We assume no prior knowledge of the initialization of the parameters. Consequently, all parameters are initialized with non-informative priors detailed below. The true number of NSDoH profiles is unknown. Using the upper bound $K_{max}$( also referred to as $K$ in our model description above), we impose an overfitted finite mixture model[45, 46]. We fit the model with a large upper bound on the number of clusters, $K=50$, coupled with a prior on the number of clusters $K<K_{max}$, where the model treats $K$ as another unknown parameter, and allow a data-driven approach to estimating $K$[45, 24, 46]. We impose the following priors on model parameters:
@@
$\boldsymbol{\pi}\mid K\sim\text{Dirichlet}(\gamma_{1},\dots,\gamma_{K}),\quad\text{where }\gamma_{k}=1\;\forall\;k.$
@@
The following full conditional distributions were used to update the model parameters:
@@
Given that they are smaller geographic units, some census tracts have missing information on some of the NSDoH variables (ranging from 14 to 24 census tracts with missing data across all variables). This model can handle missing data by imputing these values using the parameter’s posterior mean estimates. The $MC^{3}$ for posterior computation was run for 15,000 iterations, with thinning every 10 iterations (i.e., retaining every 10th sample) and the first 5000 iterations removed as part of posterior samples post-processing. For the MCMC chain heating parameter denoted as $h_{m},m=2,...,M$, where $M$ is the total number of parallel chains, we used incremental heating where the heat of the $m$’th chain is $h_{m}=1/[1+\Delta T\times(m-1)]$, and we tuned the parameter $\Delta T$ such that swaps between chains were accepted 20%- 60% of time [57]. Through this process, we found that four (4) heated chains for the $MC^{3}$ algorithm produced good mixing, and for this type of data, smaller $\Delta T=0.01$ works better. Posterior mean estimates were calculated from the remaining 1000 iterations collected from the sampler’s output. The generated MCMC samples were postprocessed using the Equivalence Classes Representatives (ECR) algorithm to overcome label-switching identifiability issues inherent in Bayesian mixture models[58]. The most probable number of clusters ($K_{map}$) given the data was inferred, and the NSDoH profile assignment probabilities for each census tract were subsequently estimated after reordering with the ECR algorithm given $K_{map}$. Table 1.1 presents the median and interquartile range (IQR) of the assignment probabilities for census tracts to their most probable profile, while Figure 1.2 illustrates the distribution of assignment probabilities across all NSDoH profiles given $K_{map}$. Some cluster assignment probabilities were relatively low; however, they still represented the highest probabilities among all cluster options for each census tract given $K_{map}$. These lower probabilities suggest the presence of unobserved heterogeneity, potentially driven by unmeasured NSDoH variables that we did not include in our model. Final clusters were qualitatively described based on thematic domains to define the NSDoH profiles, as discussed in Section 3.1 of the manuscript.
Additionally, we conducted sensitivity analyses using alternative priors (and combinations of priors) available in the software package. Specifically, we considered $K\sim Uniform(1,K_{max})$, and $\boldsymbol{\pi}|K\sim\text{Dirichlet}(\frac{1}{K_{max}},...,\frac{1}{K_{max}})$. Under these priors, we observed poorer MCMC mixing, as indicated by lower swap acceptance probabilities between heated chains. The estimated number of clusters consistently exceeded 10, and the cluster size distributions exhibited a combination of large and smaller clusters, including occasional singleton clusters. The emergence of singleton clusters is consistent with using a sparse Dirichlet prior on the mixing weights, which approximates a Dirichlet process and promotes the formation of smaller clusters.
Table 1.1: MBMM estimated assignment probability of census tracts to the most probable profile.
NSDoH Profile Median (IQR)
Profile 1 0.946 (0.161)
Profile 2 0.972 (0.0581)
Profile 3 0.923 (0.200)
Profile 4 0.874 (0.233)
Profile 5 0.865 (0.259)
Figure 1.3: Distribution of assignment probabilities for each NSDoH profile.
@@
Table 1.2 shows the distribution of age and race/ethnicity within the census tracts used to further characterize NSDoH profiles in Section 3.2 of the manuscript.
@@
Table 1.2: Distribution of population characteristics within census tracts in a given NSDoH Profile.
@@
Profile 1 Profile 2 Profile 3 Profile 4 Profile 5
Mean (SD) n = 468 n = 375 n= 263 n = 236 n = 136
Census tract population 5060 (1860) 4220 (1710) 4520 (1800) 4500 (1810) 4770 (1660)
Median Age 45.6 (6.05) 35.3 (5.29) 43.3 (6.88) 38.0 (6.98) 34.6 (5.79)
Ages 20-24 5.35 (4.94) 8.10 (5.15) 7.19 (6.98) 7.78 (5.95) 10.8 (9.50)
Ages 25-34 9.12 (3.30) 16.3 (4.66) 12.3 (3.93) 20.4 (12.1) 24.0 (9.04)
Ages 35-44 11.4 (2.88) 13.1 (3.34) 11.4 (3.18) 12.8 (3.17) 12.8 (4.06)
Age 45-54 15.4 (3.15) 12.5 (3.14) 13.6 (3.42) 12.5 (4.22) 10.7 (3.93)
Age 55-59 8.29 (2.02) 6.06 (1.93) 7.91 (2.69) 6.09 (2.44) 5.48 (2.34)
Ages 60-64 7.66 (2.30) 5.44 (2.15) 7.38 (2.40) 5.47 (2.32) 5.10 (2.20)
Ages 65-74 11.4 (4.11) 7.15 (2.68) 10.7 (3.92) 8.55 (4.93) 7.21 (3.00)
Ages 75-84 5.51 (2.57) 3.59 (1.96) 5.42 (2.52) 4.35 (2.14) 3.60 (1.88)
Age 85 or older 2.55 (1.81) 1.74 (1.30) 2.72 (1.95) 2.28 (1.93) 2.02 (1.75)
Non-Hispanic White 89.0 (7.05) 41.9 (24.9) 82.0 (14.6) 74.7 (10.5) 57.0 (13.2)
Non-Hispanic Black 1.69 (2.22) 16.2 (19.0) 4.58 (6.93) 4.30 (4.93) 11.1 (9.91)
Non-Hispanic Asian 4.07 (4.99) 5.63 (9.05) 2.75 (4.12) 11.4 (8.04) 14.3 (9.65)
Hispanic or Latino 3.14 (2.58) 31.9 (22.4) 7.61 (9.18) 6.28 (3.86) 14.0 (9.48)
  • Shorthand names for NSDoH Profiles: 1) advantaged non-Hispanic White, 2) disadvantaged racially/ethnically diverse (BHL+; non-Hispanic Black (B) and Hispanic/Latino (HL)), more renter-occupied housing with limited EN proficiency, 3) working class lower educational attainment, 4) racially/ethnically diverse (A+; non-Hispanic Asian (A)) and greater economic security and educational attainment, 5) racially/ethnically diverse (ABHL+), more renter-occupied housing with limited EN proficiency.
2 Regression Analysis Additional
The main outcome is optimal care, defined as adherence to National Comprehensive Cancer Network (NCCN) guidelines. These guidelines recommend a combination of therapies depending on the stage and grade of the tumor (Figure 2.1). We used year-specific guidelines spanning the study period due to gradual changes in NCCN treatment guidelines. We determined optimal care by comparing the treatment received and the treatment recommended by NCCN. For example, we know that surgery is the first course of treatment, and thus if a woman received surgery alone or surgery and other additional therapies, as shown in Figure 2.1, based on the stage and grade of their tumor, then she is classified as having received optimal care. We combined all the corresponding data to create a binary outcome variable, where: